Circular No. 7964

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002ev, 2002ew, 2002ex, 2002ey Four new supernovae have been discovered on unfiltered CCD images,

as listed below. SN 2002ey was found on LOTOSS KAIT images (cf. IAUC 7906) by D. Hutchings and W. D. Li, and the other three were found by W. M. Wood-Vasey et al. (cf. IAUC 7916) on Palomar NEAT images.

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  Offset
2002ev Aug. 24.26 20^{h}23^{m}00^{s}.85 − 1^{o}48′14′′.6 18.8 8′′ E, 7′′ S
2002ew Aug. 24.34 20 54 30.50 − 0 08 26.0 17.8 4′′ S
2002ex Aug. 19.28 22 09 00.79 −10 36 25.8 19.5 1′′.5 N
2002ey Sept. 3.3 23 10 12.33 + 7 32 59.9 17.7: 5′′.8 E, 5′′.5 N

Additional magnitude estimates with the same instrumentation for each respective object: SN 2002ev, June 8 UT, [20.5; July 20, [20.5; Aug. 4, [20.5; 15, 20.0–20.5 (hint); 24.31, 18.8; 25, 19.1; 31, 18.7; Sept. 3, 18.7. SN 2002ew, June 10, [19.5; Aug. 4, [19.5; 24.40, 17.8; 25, 17.7. SN 2002ex, July 22, [20.5; Aug. 8, [20.5; 17, 20.0–20.5 (hint); 19.40, 19.5; 28, 18.7. SN 2002ey, Aug. 17.3, [19.0:; Sept. 4.3, 17.7:.

R. Chornock, W. D. Li, and A. V. Filippenko, University of California at Berkeley (UCB), report that CCD spectra (range 310–1000 nm), obtained on Sept. 4 UT with the Shane 3-m reflector at Lick Observatory, reveal that SN 2002ey is a type-Ia supernova resembling the peculiar subluminous object SN 1991bg (e.g., Filippenko et al. 1992, A.J. 104, 1543). SN 2002ew is a young type-II supernova. The object has a mostly featureless blue continuum with narrow (unresolved, perhaps from an H II region) Hα emission present at a redshift of 8920 km/s. There is a very weak, lowcontrast, broader component of Hα emission, as well. SN 2002ev is also of type II, with a strong Hα line exhibiting a P-Cyg profile. SN 2002ex is a type-Ic supernova about a week past maximum brightness. Strong emission lines from a superposed H II region yield a redshift of 11310 km/s.

SUPERNOVAE 2002em AND 2002ep Chornock, B. Swift (also at UCB), Filippenko, and Li add that inspec-

tion of CCD spectra, obtained as above at Lick on Sept. 3, reveal that SN 2002ep (IAUC 7959) is of type Ia, ∼ 1 month past maximum brightness. SN 2002em (IAUC 7955) is a type-II supernova, but its spectrum consists mostly of a smooth continuum with a relatively narrow (FWHM ∼ 1500 km/s) Hα emission feature.

2002 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green